# ECORE: Energy-Conscious Optimized Routing for Deep Learning Models at the Edge

### v2

@@ 4.1.1. Dataset
We also evaluated our routing pipeline on a real-world video stream of pedestrians crossing the road by treating it as a detection dataset. Because no manual annotations were available, we decoded the raw file and extracted every frame (at the original frame rate) into a sequence of JPEG images. We then generated the ground-truth annotations by running the powerful YOLOv8x model over each frame—saving its detections in YOLO-format text files to serve as our reference labels.
Figure 3. Distribution of Object Counts per Image in the COCO Validation Set.
4.1.2. Testbed
@@
This section outlines the experimental testbed used to evaluate the proposed routing approach. The testbed configuration is informed by extensive prior benchmarking results (Alqahtani et al., 2024), which profiled various combinations of edge devices and object detection models across three key performance metrics: energy consumption, inference time, and mAP. The devices evaluated included Raspberry Pi 3 (with and without TPU), Raspberry Pi 4 (with and without TPU), Raspberry Pi 5 (with and without TPU and with an AI HAT), and the Jetson Orin Nano, making a total of eight. Eight object detection models profiled ranged from lightweight variants—SSD v1, SSD Lite, EfficientDet-Lite0, Lite1, and Lite2 to the YOLOv8 family (Nano, Small, and Medium). These metrics collectively reflect the trade-offs inherent in deploying object detection models on edge hardware. Notably, the benchmarking revealed that no single model–device pair consistently outperforms others across all criteria. Therefore, model selection must be context-dependent, guided by the specific performance objective. Figure 4 presents the accuracy–energy consumption trade-offs for all possible 64 model–device combinations.
@@
To construct the testbed, we selected only model–device pairs that lie on or near the Pareto front in at least one performance dimension. We repeated the all the configurations in Figure 4 per 5 object groups, i.e., ‘0’, ‘1’, ‘2’, ‘3’, and ‘4 or more’ objects per image. This ensures that all included devices offer meaningful trade-offs—i.e., each is strong in at least one metric—while excluding configurations that are consistently suboptimal across all criteria. Energy efficiency and latency do not vary across groups. Therefore, for energy efficiency, we selected Jetson Orin Nano running SSD v1 which demonstrated the lowest energy consumption among all configurations. In terms of inference latency, the Raspberry Pi 5 with TPU paired with SSD v1 achieved the shortest inference time. For detection accuracy (mAP), the optimal configuration varied across object count groups, underscoring the need for adaptive routing. Specifically, for Group 1 (images with no objects), the Raspberry Pi 5 with TPU running SSD v1 achieved the highest mAP. For Group 2 (images with one object), the Raspberry Pi 5 with TPU running SSD Lite performed best. In Group 3 (images with two objects), the Jetson Orin Nano with YOLOv8-small yielded the highest accuracy. Finally, for Groups 4 and 5 (images with three and four or more objects), the Raspberry Pi 5 equipped with an AI Hat running YOLOv8-small demonstrated superior performance. Table 1 provides the finalized list of device-model pairs, and Table 2 details the corresponding device specifications. This testbed supports a thorough assessment of the routing algorithm’s effectiveness in achieving optimal trade-offs between energy efficiency and performance across a variety of edge computing environments.
@@
Figure 4. Pareto frontiers comparing object detection models across various edge devices.
Table 1. Experimental Testbed Configurations
Metrics Edge Device Object Detection Model Framework
@@
Inference Time Raspberry Pi 5 SSD v1 TFLite
@@
mAP – Group 2 Raspberry Pi 5 SSD Lite TFLite
@@
Table 2. Testbed Device Specifications
@@
Raspberry Pi 5 CPU 4 GB Debian Bookworm Off Off
@@
4.2. Experimental Procedure
@@
  • Oracle Router (Orc): Serves as an ideal benchmark by assuming perfect prior knowledge of the number of objects in each incoming image. This is achieved by leveraging the ground truth annotations. The true object count is passed as metadata alongside the image request to the gateway, enabling the routing algorithm to select the optimal model–device pair by applying the $\delta$mAP criterion. This router reflects the upper bound of routing performance under perfect estimation.
@@
Given the varying performance requirements of different real-world applications, we evaluated the proposed and baseline routers across a range of $\delta$mAP values (e.g., 0, 5, 10, 15, 20, and 25). A value of $\delta$mAP = 0 enforces strict accuracy, whereas higher values allow greater flexibility in selecting less accurate but more energy-efficient models.
To ensure consistent load generation, we used Locust (Jonatan et al., 2025) to send image requests back-to-back in a piggybacked fashion—each new request is triggered only after the previous image has been processed and a response is received, simulating continuous workloads. Performance was evaluated using the following primary metrics:
@@
  • Accuracy: Evaluated using the mAP, computed via the FiftyOne tool (Voxel51, 2025), which provides robust comparisons against ground truth annotations.
@@
4.3. Experimental Results
4.3.1. Full COCO Dataset Results:
To evaluate the performance of the proposed routing approaches in comparison to baseline methods, we conducted experiments using the full COCO validation dataset, with the $\delta$ mAP threshold set to 5 unless otherwise specified. The results across energy consumption, latency, and detection accuracy are presented in Figure LABEL:fig:results_full_coco_dataset. In terms of dynamic energy consumption Figure LABEL:fig:results_full_coco_dataset(b), LE achieves the lowest energy usage, establishing the theoretical lower bound. LI follows closely, consuming only marginally more. On the other end of the spectrum, HMG, achieving best accuracy per group, increases energy consumption by approximately 80 percent over the lower bound. The proposed SF exhibits an even greater rise of nearly 85%, attributed to the additional overhead introduced by gateway analysis. The Orc router results in only 50% increase in energy usage. Among the proposed methods, ED and OB increase consumption by about 45% and 37%, respectively. In comparison, RR and Rnd lead to energy increases of around 55% and 50%.
With respect to detection accuracy (mAP), Figure LABEL:fig:results_full_coco_dataset(a) shows that HMG achieves the highest score, setting the upper bound. Orc and SF perform similarly, with accuracy dropping by less than one percent. HM and ED experience slight reductions of approximately 1% and 2%, respectively. A more significant 9% drop is observed in the OB router’s performance, attributed to the highly variable and unpredictable object counts present in the COCO dataset. Finally, RR and Rnd incur significant drops of around 25%, while LE and LI show the largest losses—about 40% and 50%, respectively. Results show that the proposed routing approaches, such as ED and OB, can achieve mAP values very close to optimal methods like HMG, while also delivering significant energy savings of approximately 45% and 37%, respectively.
Latency results, measured as the total time to process the entire dataset, are presented in Figure LABEL:fig:results_full_coco_dataset(c). As expected, LI achieves the lowest overall latency, with LE following closely—incurring an increase of less than 1%. In contrast, HM and SF exhibit the highest latency increases, each rising by approximately 75%, while HMG incurs a latency overhead of around 60%.. In contrast, Orc, OB, and ED introduce relatively small latency increases of about 11%.
@@ 4.3.1. Full COCO Dataset Results:
4.3.2. Balanced Sorted Dataset Results
This section highlights the results of evaluating our proposed routing approaches and baseline routing strategies using the balanced sorted dataset as shown in Figure LABEL:fig:results_sorted_balanced_dataset. With respect to dynamic energy consumption Figure LABEL:fig:results_sorted_balanced_dataset(b), LE demonstrates the most energy-efficient performance, consuming only 227 mWh and establishing the theoretical lower bound. LI performs similarly, with only a slight increase in consumption. At the opposite end, HMG exhibits a substantial rise in energy usage—approximately 50% higher than the baseline—while SF records more than double the energy consumption of LE. The Orc router shows a 40% increase in energy use relative to the lower bound. Among the proposed approaches, ED and OB consume 64% and 45% more energy, respectively. In comparison, the baseline methods RR and Rnd result in increases of roughly 70% and 62%.
Latency results, presented in Figure LABEL:fig:results_sorted_balanced_dataset(c), show that LI achieves the lowest total latency at approximately 306 seconds, with LE following closely behind, incurring only a 2-second increase. HM and SF exhibit the highest latency overheads, with increases of around 77% and 81%, respectively. HMG also shows a notable rise, adding approximately 34% to the baseline latency. In contrast, Orc and OB introduce modest increases of about 9%, while ED shows a slightly higher increase of 13%. The RR and Rnd routers fall in the mid-range, with latency increases of 18% and 22%, respectively.
In terms of detection accuracy (mAP), Figure LABEL:fig:results_sorted_balanced_dataset(a) indicates that HMG achieves the highest score at 40.94, establishing the upper performance bound. Orc, SF, and OB closely follow, each exhibiting a negligible decline of less than 1%. ED also shows minor reductions of approximately 1%. In contrast, RR and Rnd suffer notable accuracy losses of around 18%, while LE and LI demonstrate the most substantial declines, with reductions of approximately 30% and 40%, respectively.
@@ 4.3.2. Balanced Sorted Dataset Results
4.3.3. Video Dataset Results
This section presents the results of the different routing approaches evaluated using the pedestrian video dataset. The corresponding Figure LABEL:fig:results_video_dataset illustrates the performance trends across key metrics. With respect to dynamic energy consumption Figure LABEL:fig:results_video_dataset(b), LE consumes only 85 mWh. LI follows with a 75% increase in consumption over LE. At the opposite end, HMG shows a substantial rise in energy usage—approximately 100% higher than the baseline—while SF consumes more than three times the energy of LE, exceeding 200%. The Orc router results in an 85% increase relative to the lower bound. Among the proposed methods, ED and OB increase energy usage by approximately 138% and 100%, respectively. For comparison, the baseline methods RR and Rnd show similar trends, each exceeding 200% of the energy consumption of LE.
Latency results, illustrated in Figure LABEL:fig:results_video_dataset(c), indicate that LI achieves the best performance, with the lowest total latency of approximately 164 seconds. LE follows closely, incurring a minimal increase of 7 seconds. On the higher end, HM and SF experience the greatest latency overheads, with increases of approximately 61% and 73%, respectively. HMG also shows a significant rise, adding around 50% over the baseline. In comparison, Orc and OB introduce only modest latency increases of about 9%, while ED results in a more pronounced increase of 40%. The RR and Rnd routers occupy a middle ground, with latency overheads of 29% and 17%, respectively.
Regrading accuracy (mAP), Figure LABEL:fig:results_video_dataset(a) shows that HMG achieves the highest score at 79.95, setting the upper bound for performance. Orc closely follows, with a minimal decline of less than 1%. SF, OB, and HM each experience moderate reductions of around 4%. ED shows a more significant drop, with accuracy decreasing by approximately 14%. In contrast, RR and Rnd incur substantial losses of 50% and 55%, respectively, while LE and LI show the most severe accuracy degradation, with mAP reductions of about 63% and 75%.
@@ 4.3.3. Video Dataset Results
4.3.4. Results Across Different $\delta$mAP Values
This section presents the performance of the Oracle and proposed routers across varying $\delta$mAP values. All routers including the Oracle show a substantial reduction in energy consumption when $\delta$mAP increases from 0 to 5. Between $\delta$mAP values of 5 and 20, energy consumption continues to decrease slightly, followed by a more noticeable drop at $\delta=25$ as Figure LABEL:fig:results_differnt_delta depicts.
A similar trend is observed for total latency. All routers exhibit a sharp decline in latency when $\delta$mAP increases from 0 to 5, with only marginal improvements observed between values of 5 and 25.
In terms of detection accuracy (mAP), most routers maintain stable performance as $\delta$mAP increases from 0 to 5. However, a slight decrease in mAP is observed between $\delta=5$ and $\delta=15$, followed by a more significant decline between $\delta=20$ and $\delta=25$.
Insight #4. Results show that even setting $\delta$mAP to a small value such as 5—allowing up to a 5% tolerance in accuracy drop—yields significant reductions in both energy consumption and latency, while maintaining detection performance comparable to the strictest setting ($\delta=0$). Notably, with $\delta$mAP set to 5, the actual accuracy drop remains relatively low, around 2%.
4.4. Discussions
@@
5. Related Work
Table 3. Comparison of Related Work and ECORE
@@ 5. Related Work
Ji et al. (Ji et al., 2025)
Trinh et al. (Trinh et al., 2018)
Tu et al. (Tu et al., 2023)
Zhang et al. (Zhang et al., 2022)
Tundo et al. (Tundo et al., 2023)
Matathammal et al. (Matathammal et al., 2025)
Kulkarni et al. (Kulkarni et al., 2023)
Marda et al. (Marda et al., 2024)
Stripelis et al. (Stripelis et al., 2024)
Maurya et al. (Maurya et al., 2024)
Zheng et al. (Zheng et al., 2025)
Guha et al. (Guha et al., 2024)
Mohammadshahi et al. (Mohammadshahi et al., 2024)
Sikeridis et al. (Sikeridis et al., 2024)
@@
This section discusses related studies and compares them with our proposed work as Table 3 depicts. Ji et al. (Ji et al., 2025) introduce a Cloud-Edge Large-Tiny Model Collaborative (CELTC) architecture that employs deep reinforcement learning to enhance surveillance data processing and decision-making in offshore drilling platform environments. The CELTC framework integrates edge–cloud computing by deploying high-accuracy large models on cloud servers and lightweight, efficient models on edge devices to balance computational demand and responsiveness. Similarly, Trinh et al. (Trinh et al., 2018) explore the use of Mobile Edge Computing (MEC) to address energy management challenges in power-constrained IoT devices while enabling low-latency processing of high-resolution visual data. Focusing on facial recognition for disaster response scenarios, the authors propose an offload decision-making algorithm that dynamically determines whether to process tasks at the edge or offload them to the cloud, based on current workload and energy constraints. While these studies present valuable edge–cloud collaborative architectures, their focus differs from ours. Specifically, they rely on cloud resources to support edge processing, whereas our work emphasizes fully distributed collaboration among multiple edge devices. Furthermore, our approach directly targets the optimization of energy consumption and latency in edge environments without compromising detection accuracy, making it well-suited for real-time applications where both efficiency and performance are critical.
Tu et al. (Tu et al., 2023) propose EdgeLeague, a framework designed to manage multiple video streams with diverse Quality of Service (QoS) requirements, ensuring reliable surveillance performance under constrained edge resources and fluctuating uplink bandwidth. EdgeLeague leverages collaborative edge computing and optimizes camera network configurations to solve an NP-hard integer nonlinear problem by dynamically adjusting camera resolutions and selecting appropriate detection models across cooperating edge nodes. Similarly, Zhang et al. (Zhang et al., 2022) address configuration selection and bandwidth allocation for multiple video streams connected to a shared uplink channel through a common edge node. The authors propose JCAB, an efficient online algorithm that jointly optimizes configuration adaptation and bandwidth distribution. The approach is further extended to a multi-edge scenario, allowing individual users or video streams to select their preferred edge node, thereby enhancing flexibility and resource utilization. While these studies focus on optimizing system performance by adapting hardware configurations and bandwidth allocation, our approach takes a different direction. Instead of adjusting device settings, we analyze the content of incoming images and dynamically route requests to the most suitable edge device–model pair. This enables energy-efficient inference while preserving detection accuracy, offering a lightweight and scalable alternative to hardware-level reconfiguration strategies.
Tundo et al. (Tundo et al., 2023) introduces an energy-conscious strategy for designing and deploying self-adaptive, AI-driven applications, aiming to balance application-level objectives—such as object detection accuracy and frame processing rate—with energy consumption. The authors address the challenge of identifying optimal system configurations for self-adaptation using a meta-heuristic search method that requires minimal empirical sampling. Similarly, Matathammal et al. (Matathammal et al., 2025) present EdgeMLBalancer, a self-adaptive methodology designed to optimize CPU utilization and resource management on edge devices. Their approach dynamically switches between deployed models based on real-time CPU usage monitoring across processor cores. In Kulkarni et al. (Kulkarni et al., 2023), the concept of a Machine Learning Model Balancer is proposed to manage uncertainty in ML model outputs through the coordinated use of multiple models. Building on this idea, the authors introduce AdaMLS, which enhances the traditional MAPE-K (Monitor-Analyze-Plan-Execute over a Knowledge base) loop to enable continuous adaptation of machine learning systems. The study in Marda et al. (Marda et al., 2024) introduces SWITCH, a framework designed to improve system-level adaptability by enabling runtime model switching. SWITCH functions as a general-purpose web service applicable to various ML scenarios, and is specifically demonstrated in an object detection use case. Whereas these studies focus on self-adaptive techniques that switch between models within a single edge device, our work adopts a broader architectural perspective. We design a system that maintains a pool of edge devices and model variants, and introduces a dynamic routing mechanism to select the most appropriate device–model pair for each image. This approach supports distributed, energy-efficient processing across multiple edge nodes while preserving detection accuracy, extending adaptability beyond local device-level switching to collaborative edge inference.
Recent research has explored various routing strategies to improve the efficiency and adaptability of Large Language Model (LLM) inference. Stripelis et al. (Stripelis et al., 2024) introduce a modular querying system that dynamically dispatches incoming queries to specialized LLM experts based on the query’s requirements. Similarly, Maurya et al. (Maurya et al., 2024) propose an algorithm that selects the most appropriate subset of LLMs from a larger model pool to optimize performance and cost. Zheng et al. (Zheng et al., 2025) present CITER, a fine-grained token-level routing framework that enables collaborative inference between small and large language models by directing low-importance tokens to smaller models for efficiency and reserving complex tokens for larger models to maintain accuracy. Guha et al. (Guha et al., 2024) introduce Smoothie, a routing method based on weak supervision that constructs a latent variable graphical model from the embeddings of multiple LLM responses, enabling accurate output estimation without labeled data. Mohammadshahi et al. (Mohammadshahi et al., 2024) offers Routoo, a flexible routing architecture that balances latency, cost, and inference quality by allowing users to control trade-offs according to task requirements. Sikeridis et al. (Sikeridis et al., 2024) propose PickLLM which employs a reinforcement learning–based framework that dynamically assigns queries to models using a reward function guided by query-specific metrics such as accuracy, latency, and cost. While these works demonstrate significant advances in LLM routing, they are primarily focused on natural language processing tasks in cloud-based environments. In contrast, our work addresses object detection in edge computing settings, with a particular emphasis on minimizing energy consumption and processing latency and accuracy factors that are critical for real-time, resource-constrained applications.
6. Conclusions and Future Direction
Deploying object detection models on resource-constrained edge devices—particularly those powered by batteries or sustainable energy sources such as solar panels—requires careful consideration of energy consumption. In many real-world applications, especially safety-critical systems, the operating environment is highly dynamic, and the number of objects captured in each image can vary significantly. This variability can substantially affect the accuracy of object detection models, necessitating adaptive strategies that balance efficiency and performance.
To address these challenges, we proposed ECORE along with three dynamic, estimation-based routing algorithms that first estimate the number of objects in an image using lightweight pre-processing techniques. Then, based on this estimation, the algorithms route the request to the most suitable edge model-device pair. The routing decision is informed by profiling and is designed to minimize overall energy consumption while preserving detection accuracy within acceptable range.
Our experimental evaluation demonstrates that the proposed routing strategy performs favorably compared to baseline approaches. Specifically, it achieves a substantial reduction in energy usage while maintaining accuracy levels within an acceptable margin. The results highlight the potential of dynamic, context-aware routing for improving the energy-efficiency of AI workloads at the edge.

### v3

@@ IV-A1 Dataset
We also evaluated our routing pipeline on a real-world video stream of pedestrians crossing the road by treating it as a detection dataset. Because no manual annotations were available, we decoded the raw file and extracted every frame (at the original frame rate) into a sequence of JPEG images. We then generated the ground-truth annotations by running the powerful YOLOv8x model over each frame, saving its detections in YOLO-format to serve as our reference labels.
Figure 3: Distribution of Object Counts / Image in COCO dataset.
IV-A2 Testbed
@@
This section outlines the experimental testbed used to evaluate the proposed routing approach. The testbed configuration builds on exhaustive prior benchmarking results [5], which involved a comprehensive and resource-intensive evaluation of various combinations of edge devices and object detection models across three key performance metrics: energy consumption, inference time, and mAP. This benchmarking effort, which profiled a wide range of devices and models, provides an invaluable reference point for the research community studying edge-based object detection. Specifically, the devices evaluated included Raspberry Pi 3 (with and without TPU), Raspberry Pi 4 (with and without TPU), Raspberry Pi 5 (with and without TPU and with an AI HAT), and the Jetson Orin Nano, making a total of eight. The eight profiled object detection models spanned from lightweight variants—SSD v1, SSD Lite, EfficientDet-Lite0, Lite1, and Lite2—to the YOLOv8 family (Nano, Small, and Medium). Together, these metrics reflect the inherent trade-offs in deploying object detection models on edge hardware. Notably, the benchmarking revealed that no single model–device pair consistently outperforms others across all criteria, reinforcing the need for context-dependent model selection guided by specific performance objectives. Fig. 4 illustrates the accuracy–energy consumption trade-offs for all 64 model–device combinations.
@@
To construct our testbed, we selected only model–device pairs that lie on or near the Pareto front in at least one performance dimension. We repeated all the configurations in Fig. 4 for five object groups—‘0’, ‘1’, ‘2’, ‘3’, and ‘4 or more’ objects per image—an effort that required extensive experimentation and significant computational time. This ensures that all included devices offer meaningful trade-offs—i.e., each is strong in at least one metric—while excluding configurations that are consistently suboptimal across all criteria. Energy efficiency and latency do not vary across groups. Therefore, for energy efficiency, we selected Jetson Orin Nano running SSD v1 which demonstrated the lowest energy consumption among all configurations. In terms of inference latency, the Raspberry Pi 5 with TPU paired with SSD v1 achieved the shortest inference time. For detection accuracy (mAP), the optimal configuration varied across object count groups, underscoring the need for adaptive routing. Specifically, for Group 1 (images with no objects), the Raspberry Pi 5 with TPU running SSD v1 achieved the highest mAP. For Group 2 (images with one object), the Raspberry Pi 5 with TPU running SSD Lite performed best. In Group 3 (images with two objects), the Jetson Orin Nano with YOLOv8-small yielded the highest accuracy. Finally, for Groups 4 and 5 (images with three and four or more objects), the Raspberry Pi 5 equipped with an AI Hat running YOLOv8-small demonstrated superior performance. Table I provides the finalized list of device-model pairs, and Table II details the corresponding device specifications. This testbed supports a thorough assessment of the routing algorithm’s effectiveness in achieving optimal trade-offs between energy efficiency and performance across a variety of edge computing environments.
@@
Figure 4: Pareto frontiers comparing object detection models across various edge devices.
TABLE I: Experimental Testbed Configurations
Metrics Edge Device Object Detection Model Runtime
@@
Inference Time Raspberry Pi 5 + TPU SSD v1 TFLite
@@
mAP – Group 2 Raspberry Pi 5 + TPU SSD Lite TFLite
@@
TABLE II: Testbed Device Specifications
@@
Pi 5 + Coral TPU TPU 4 GB Debian Bookworm Off Off
@@
Jetson Orin Nano GPU 8 GB JetPack 5.1.3 Off Off
IV-B Experimental Procedure
@@
[…]
@@
[…]
To ensure consistent load generation, we used Locust  [^1] https://locust.io/ to send image requests back-to-back in a piggybacked fashion where each new request is triggered only after the previous image has been processed and a response is received, simulating continuous workloads. Performance was evaluated using the following primary metrics:
@@
  • Accuracy: Evaluated using the mAP, computed via the FiftyOne tool  [^2] https://docs.voxel51.com/, which provides robust comparisons against ground truth annotations.
@@
IV-C Experimental Results
IV-C1 Full COCO Dataset Results:
To evaluate the performance of the proposed routing approaches in comparison to baseline methods, we conducted experiments using the full COCO validation dataset, with the $\delta$ mAP threshold set to 5 unless otherwise specified. The results across energy consumption, latency, and detection accuracy are presented in Fig. 7(a). In terms of dynamic energy consumption Fig. 7(a)(b), LE achieves the lowest energy usage, establishing the theoretical lower bound. LI follows closely, consuming only marginally more. On the other end of the spectrum, HMG, achieving best accuracy per group, increases energy consumption by approximately 80 percent over the lower bound. The proposed SF exhibits an even greater rise of nearly 85%, attributed to the additional overhead introduced by gateway analysis. The Orc router results in only 50% increase in energy usage. Among the proposed methods, ED and OB increase consumption by about 45% and 37%, respectively. In comparison, RR and Rnd lead to energy increases of around 55% and 50%.
With respect to detection accuracy (mAP), Fig. 7(a)(a) shows that HMG achieves the highest score, setting the upper bound. Orc and SF perform similarly, with accuracy dropping by less than one percent. HM and ED experience slight reductions of approximately 1% and 2%, respectively. A more significant 9% drop is observed in the OB router’s performance, attributed to the highly variable and unpredictable object counts present in the COCO dataset. Finally, RR and Rnd incur significant drops of around 25%, while LE and LI show the largest losses—about 40% and 50%, respectively. Results show that the proposed routing approaches, such as ED and OB, can achieve mAP values very close to optimal methods like HMG, while also delivering significant energy savings of approximately 45% and 37%, respectively.
Latency results, measured as the total time to process the entire dataset, are presented in Fig. 7(a)(c). As expected, LI achieves the lowest overall latency, with LE following closely—incurring an increase of less than 1%. In contrast, HM and SF exhibit the highest latency increases, each rising by approximately 75%, while HMG incurs a latency overhead of around 60%.. In contrast, Orc, OB, and ED introduce relatively small latency increases of about 11%.
@@ IV-C1 Full COCO Dataset Results:
[…]
IV-C2 Balanced Sorted Dataset Results
This section highlights the results of evaluating our proposed routing approaches and baseline routing strategies using the balanced sorted dataset as shown in Fig. 7(b). With respect to dynamic energy consumption Fig. 7(b)(b), LE demonstrates the most energy-efficient performance, consuming only 227 mWh and establishing the theoretical lower bound. LI performs similarly, with only a slight increase in consumption. At the opposite end, HMG exhibits a substantial rise in energy usage—approximately 50% higher than the baseline—while SF records more than double the energy consumption of LE. The Orc router shows a 40% increase in energy use relative to the lower bound. Among the proposed approaches, ED and OB consume 64% and 45% more energy, respectively. In comparison, the baseline methods RR and Rnd result in increases of roughly 70% and 62%.
Latency results, presented in Fig. 7(b)(c), show that LI achieves the lowest total latency at approximately 306 seconds, with LE following closely behind, incurring only a 2-second increase. HM and SF exhibit the highest latency overheads, with increases of around 77% and 81%, respectively. HMG also shows a notable rise, adding approximately 34% to the baseline latency. In contrast, Orc and OB introduce modest increases of about 9%, while ED shows a slightly higher increase of 13%. The RR and Rnd routers fall in the mid-range, with latency increases of 18% and 22%, respectively.
In terms of detection accuracy (mAP), Fig. 7(b)(a) indicates that HMG achieves the highest score at 40.94, establishing the upper performance bound. Orc, SF, and OB closely follow, each exhibiting a negligible decline of less than 1%. ED also shows minor reductions of approximately 1%. In contrast, RR and Rnd suffer notable accuracy losses of around 18%, while LE and LI demonstrate the most substantial declines, with reductions of approximately 30% and 40%, respectively.
@@ IV-C2 Balanced Sorted Dataset Results
[…]
IV-C3 Video Dataset Results
This section presents the results of the different routing approaches evaluated using the pedestrian video dataset. The corresponding Fig. 7(c) illustrates the performance trends across key metrics. With respect to dynamic energy consumption Fig. 7(c)(b), LE consumes only 85 mWh. LI follows with a 75% increase in consumption over LE. At the opposite end, HMG shows a substantial rise in energy usage—approximately 100% higher than the baseline—while SF consumes more than three times the energy of LE, exceeding 200%. The Orc router results in an 85% increase relative to the lower bound. Among the proposed methods, ED and OB increase energy usage by approximately 138% and 100%, respectively. For comparison, the baseline methods RR and Rnd show similar trends, each exceeding 200% of the energy consumption of LE.
Latency results, illustrated in Fig. 7(c)(c), indicate that LI achieves the best performance, with the lowest total latency of approximately 164 seconds. LE follows closely, incurring a minimal increase of 7 seconds. On the higher end, HM and SF experience the greatest latency overheads, with increases of approximately 61% and 73%, respectively. HMG also shows a significant rise, adding around 50% over the baseline. In comparison, Orc and OB introduce only modest latency increases of about 9%, while ED results in a more pronounced increase of 40%. The RR and Rnd routers occupy a middle ground, with latency overheads of 29% and 17%, respectively.
Regrading accuracy (mAP), Fig. 7(c)(a) shows that HMG achieves the highest score at 79.95, setting the upper bound for performance. Orc closely follows, with a minimal decline of less than 1%. SF, OB, and HM each experience moderate reductions of around 4%. ED shows a more significant drop, with accuracy decreasing by approximately 14%. In contrast, RR and Rnd incur substantial losses of 50% and 55%, respectively, while LE and LI show the most severe accuracy degradation, with mAP reductions of about 63% and 75%.
@@ IV-C3 Video Dataset Results
[…]
(d) Accuracy, Latency and Energy of the Oracle router and proposed routers for across different delta.
IV-D Discussions
@@
V Related Work
TABLE III: Comparison of Related Work and ECORE
@@ V Related Work
Ji et al. [8]
Trinh et al. [9]
Tu et al. [10]
Zhang et al. [11]
Tundo et al. [12]
Matathammal et al. [13]
Kulkarni et al. [14]
Marda et al. [15]
Stripelis et al. [16]
Maurya et al. [17]
Zheng et al. [18]
Guha et al. [19]
Mohammadshahi et al. [20]
Sikeridis et al. [21]
Jiang et al. [22]
Kang et al. [23]
Davis and Belviranli [24]
Zhang et al. [25]
@@
This section surveys related studies and situates our approach (see Table III). We organize prior work into three scopes: (i) dynamic on-device adaptation, (ii) system-level orchestration across edge clusters and between edge and cloud, and (iii) routing policies. The first includes pipeline reconfiguration, local model and accelerator selection, and contextual-bandit policies within a single device. The second covers multi-stream configuration, bandwidth allocation, cluster scheduling, and offloading across edge and cloud. The third reviews query-aware dispatch among model pools in the cloud, which is conceptually related but outside our edge-vision scope.
V-A Dynamic on-device adaptation
Local methods adapt inference on a single device. Pipeline and content-adaptive systems such as Chameleon [22] and NoScope [23] reconfigure analytics, use cascades, or apply difference detectors to reduce cost while maintaining accuracy, all within one host. Self-adaptive switching (EdgeMLBalancer [13], SWITCH [15], AdaMLS [14]), energy-aware design (Tundo et al. [12]), and intra-chip selection among detectors and accelerators (SHIFT [24]) likewise operate within one device. On the other hand, these works optimize within a single device or processing pipeline rather than across heterogeneous edge devices.
V-B System-level orchestration
A second line of research focuses on coordinating resources across different nodes and tiers. EdgeLeague [10] and JCAB [11] adapt detector configurations and uplink bandwidth for multi-stream analytics under fluctuating network conditions, scheduling resources to meet quality-of-service and latency requirements. VideoStorm [25] performs cluster-scale scheduling based on resource and quality profiles, emphasizing throughput and deadline compliance rather than per-image content decisions. In cloud–edge offloading, CELTC [8] coordinates large cloud models with lightweight edge models using deep reinforcement learning, while Trinh et al. [9] determine when to offload vision tasks under energy constraints. These orchestration approaches optimize configurations, bandwidth, and scheduling across the system, whereas our work operates entirely at the edge and makes content-aware selections of the most suitable device–model pair for each image.
V-C Routing policies
Outside edge vision, several studies in cloud computing explore routing among large language models using techniques such as expert dispatch, subset selection, token-level collaboration, weak supervision routing, and reinforcement learning policies [16, 17, 18, 19, 20, 21]. These works share the routing concept but focus on cloud-based NLP workloads, aiming to balance quality, cost, and latency, which differs from the energy-constrained and heterogeneous nature of edge vision.
Across these threads, most prior work either depends on the cloud or optimizes configuration, bandwidth, and scheduling at the system level; adapts within a single device or pipeline; or studies routing in cloud model pools outside edge vision. In contrast, ECORE operates across edge nodes and performs content-aware routing for each image to a heterogeneous device and model pair using lightweight estimators and reducing energy while preserving detection accuracy.
VI Conclusions and Future Direction
Deploying object detection models on resource-constrained edge devices, especially those powered by batteries or renewable energy sources such as solar panels, demands careful consideration of energy consumption. In many real-world applications, particularly safety-critical systems, the environment is highly dynamic and the number of objects captured in each frame can vary greatly. This variability directly impacts the accuracy of object detection models and highlights the need for adaptive strategies that maintain a balance between efficiency and performance. To address these challenges, we introduce ECORE together with three dynamic, estimation-based routing algorithms. Each algorithm begins by applying lightweight preprocessing to estimate the number of objects in an incoming image request. Based on this estimate, the request is then routed to the most suitable edge model–device pair, a decision informed by detailed profiling of accuracy and energy trade-offs. This design aims to minimize overall energy consumption while ensuring that detection accuracy remains within an acceptable range. Through extensive experimental evaluation, we show that the proposed strategy consistently outperforms baseline approaches. In particular, it achieves significant reductions in energy usage without compromising accuracy, thereby demonstrating the effectiveness of dynamic, context-aware routing as a promising direction for enabling energy-efficient AI workloads at the edge.